The Gigawatt Power Company charges $0.10 per kilowatt-hour (kWh) for the first 1000 kWh of power purchased by a residential customer each month, but only $0.05/kWh for all additional kWh. For a residential customer with a monthly income of $400, graph the budget constraint for electric power and the composite good. The composite good has price of $1.
Answer #1

Let X= power, Y : composite good

Then if X<= 1000

BC : .1X+ Y = 400, slope = .1

If X = 1000, total spending on X = .1*1000 = 100

& With X = 1000, Y = 300

If X > 1000,

BC : .05X + Y = 400, slope = .05

So maximum possible X = 1000 + 300/.05

( As after X= 1000, Income left = 300)

= 1000+ 6000

= 7000

• BC has kink at (1000, 300)